# Remarks on classical pseudo-electrodynamics

### v1

@@ VI Conclusions
The classical pseudo-electrodynamics (PED) is a non-local electrodynamics that emerges from the dimensional reduction of the Maxwell ED, when the classical sources are constrained on a spatial plane. Consequently, this non-local ED lives in a $1+2$ space-time, and the correspondent gauge field is the candidate to mediate interactions between fermions in bidimensional materials. Thus, many applications in condensed matter physics can be investigated taking this theory as the electromagnetic sector in weyl semimetals, graphene, and others. We examine the classical aspects of the pseudo-electrodynamics (PED), as the conservation laws in the complete form, and also interesting results of the propagation fields in the radiation zone. We apply the infinitesimal symmetries in the action to obtain the Noether current of the theory. As we expect, the Noether current is not gauge invariant, and posteriorly, we introduce the technic of the Belifante-Rosenfeld tensor. Thereby, the final result is the non-local energy-momentum tensor, that is symmetric and gauge invariant. From this result, the ${\cal Z}$-component for the general angular momentum is calculated.
From the solutions for the retarded potentials, and the Jefimenko’s equations, we explore the solutions of the EM fields for an electric dipole variable with the time in the radiation zone. We calculate the radiation power of an electric dipole with constant acceleration in which the variation of the power over the radial direction is proportional to squared root of acceleration times the radial distance. This result does not reproduce the Larmor’s formula due to non-locallity of the Poynting vector in PED.

### v2

@@ VI Conclusions
The classical pseudo-electrodynamics (PED) is a non-local electrodynamics that emerges from the dimensional reduction of the Maxwell ED, when the classical sources are constrained on a spatial plane. Consequently, this non-local ED lives in a $1+2$ space-time, and the correspondent gauge field is the candidate to mediate interactions between fermions in bidimensional materials. Thus, many applications in condensed matter physics can be investigated taking this theory as the electromagnetic sector in weyl semimetals, graphene, and others. We examine the classical aspects of the pseudo-electrodynamics (PED), as the conservation laws in the complete form, and also interesting results of the propagation fields in the radiation zone. We apply the infinitesimal symmetries in the action to obtain the Noether current of the theory. As we expect, the Noether current is not gauge invariant, and posteriorly, we introduce the technics of the Belifante-Rosenfeld tensor. Thereby, the final result is the non-local energy-momentum tensor, that is symmetric and gauge invariant. From this result, the ${\cal Z}$-component for the general angular momentum is calculated.
From the solutions for the retarded potentials, and the Jefimenko’s equations, we explore the solutions of the EM fields for an electric dipole variable with the time in the radiation zone. We calculate the radiation power of an electric dipole with constant acceleration in which the variation of the power over the radial direction is proportional to squared root of acceleration times the radial distance. This result does not reproduce the Larmor’s formula due to non-locality of the Poynting vector in the PED.